Resistances of a wire is 2 ohm. It is drawn in such a way that it experiences a longitudinal strain of 200+ its new resistances is 

Grade:12

1 Answers

Arun
5 years ago
Dear student
 

R = 2 ohm

L’ = 200 %

L’ = L + (200/100)L

L’ = L + 2L = 3L

(L’/L) = 3

Vol. remains same,

A’L’ = AL

(A’/A) = (L/L’) => 1/3

A’ = A/3

R’ = p(L’/A’) = p(3L)3/A = [pL/A] 9 = R9 = 9R = 9 * 2 = 18 ohm

Regards

Arun (askIITians forum expert)